Grade 12th PassAlgebra

If one root of ax2+bx+c=0 is 3-4i and a,b,c are real,a+b+c=

ASh
8 Years agoGrade 12th Pass

1 Answer

lakshya pokharna
8 Years ago
ax^2 + bx + c=0¥;₩ be the roots of equations now ¥=3-4iand ₩=3+4i ie conjugate of ¥so ¥+₩=6¥₩=25so new equation is x^2-(¥+₩)x+¥₩=x^2-6x+25so compare coeffecients of both equation a=1;b=-6;c=25so a+b+c=20